OpenStudy (anonymous):

1. Choose the point-slope form of the equation below that represents the line that passes through the point (6, −3) and has a slope of . y − 6 = (x + 3) y = x − 6 y + 3 = (x − 6) x − 2y = 12

OpenStudy (anonymous):

y+3=(x-6) because yolo

OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the points (−3, 2) and (2, 1). y + 3 = −5(x − 2) y − 2 = −5(x + 3) y + 3 = − (x − 2) y − 2 = − (x + 3)